# A preconditioned difference of convex functions algorithm with extrapolation and line search

### v1

@@ Abstract
This paper proposes a novel proximal difference-of-convex (DC) algorithm enhanced with extrapolation and aggressive non-monotone line search for solving non-convex optimization problems. We introduce an adaptive conservative update strategy of the extrapolation parameter determined by a computationally efficient non-monotone line search. The core of our algorithm is to unite the update of the extrapolation parameter with the step size of the non-monotone line search interactively. The global convergence of the two proposed algorithms is established through the Kurdyka-Łojasiewicz properties, ensuring convergence within a preconditioned framework for linear equations. Numerical experiments on two general non-convex problems: SCAD-penalized binary classification and graph-based Ginzburg-Landau image segmentation models, demonstrate the proposed method’s high efficiency compared to existing DC algorithms both in convergence rate and solution accuracy.
@@ 1 Introduction
Here $g_{1}\left(x\right)$ and $g_{2}\left(x\right)$ are proper closed convex functions, $\nabla f\left(x\right)$ is Lipschitz continuous with constants $L$, and $X$ is a finite-dimensional Hilbert space. Such non-convex optimization formulations find broad applications in digital image processing, statistical machine learning, and related fields [1, 8, 11]. The proximal DC algorithm (DCA) with extrapolation (shortened as pDCAe henceforth) has emerged as a particularly efficient approach, successfully accelerating conventional proximal gradient methods [18, 17, 19]. This method combines the computational efficiency of proximal gradient algorithms with DC decomposition techniques, as detailed in [1, 11] for various non-convex optimizations. Our work builds upon these foundations while introducing critical enhancements to the algorithmic framework.
@@
As shown in the survey [19, Section 5], the convergence speed of DCA can be accelerated through two distinct approaches: extrapolation techniques with traditional parameterization strategies, including FISTA-based extrapolation [25, Algorithm 1] or FISTA-based restart mechanisms, e.g., [19, Algorithm 5], [29, Algorithm (3.4)] and Armijo line search methods based on energy decreasing [2, Algorithm 2], [27].
@@
Compared to the existing works, our contributions are as follows. The main contributions are a concise, lightweight non-monotone line search mechanism and a new adaptive update strategy for the extrapolation parameter. The line search completely determines the extrapolation parameter. Inspired by [2, 3, 15], we employ the aggressive non-monotone line search, where the starting point for line search is generated by the preconditioned and extrapolated DCA, instead of using the previous iteration [2, 3]. The aggressive non-monotone line search [15, Example 4] can not only relax stringent descent requirements, but can also obtain an extra step. The proposed framework enhances pDCAe by integrating these two accelerating methods. Specifically, the non-monotone line search strategy provides global guidance for local extrapolation parameter update. We refer to [31, NLSA] [23, Algorithm 1] [12, Section 4] for the traditional non-monotone line search strategy. Moreover, given the limited existing work that combines these approaches, our comparative analysis demonstrates distinct advantages over existing methods. Compared to adaptive linear search methods [30, Algorithm 1] that require the solving of auxiliary subproblems, our approach eliminates computational overhead while preserving the core efficiency of pDCAe.
@@
Furthermore, we introduce a line search framework for the case $g_{1}$ being non-smooth under new and mild conditions with a global convergence guarantee, which is not covered by the line search framework [3]. In [3], only the case $g_{2}$ is non-smooth, and $g_{1}$ with Lipschitz gradient is discussed to ensure the existence of descent directions.
@@
Numerical validations confirm the improvements. In Section 5 for numerical experiments, we compare our algorithm with the algorithms that use only a single type of acceleration method. The experiments show that our algorithm has better performance. First, for iterative sequences near critical points, the integrated update strategy stabilizes convergence by controlling the extrapolation parameter $\beta$, effectively addressing the performance degradation observed in the original implementations of pDCAe. Second, the interactive combination of extrapolation and line search accelerates convergence compared to standard extrapolation and restart approaches. These empirical results, detailed in Section 5, verify that our framework maintains the computational simplicity of pDCAe while achieving better convergence across various non-convex optimization benchmarks.
@@ 2 Notation and preliminaries
[…]
@@
[…]
@@ 3 Preconditioned DCA with extrapolation determined by line search
Now let us turn to preconditioned DCA with extrapolation for (1.1). By solving the subproblem involving the difference of convex functions by incorporating a proximal term to update $\bar{x}^{n}$, the optimization task for determining $\bar{x}^{n}$ is as follows
@@
Henceforth, we assume that the symmetric positive definite matrix $M$ serves as a preconditioning weight operator, where $\langle a,b\rangle_{M}=\langle a,Mb\rangle_{M}$ for $\forall$ $a,b\in X$ and the induced norm as $\left\|x\right\|_{M}^{2}:=\left<Mx,x\right>$. Applying first-order optimality conditions to the minimization problem (3.1) yields the critical equation for determining $\bar{x}^{n}$
$-\xi^{n}+M\left(y-y^{n}\right)+\nabla f\left(y\right)+\partial g_{1}\left(y\right)=0.$ (3.2)
@@
If $g_{1}+f$ is a quadratic function, the update $\bar{x}^{n}$ in (3.3) can be reformulated as the classical efficient preconditioned iteration, which can circumvent solving the linear equation even with mild accuracy, especially for a large-scale linear system. We will explain the preconditioning technique in Proposition 2 later.
Within the widely employed gradient descent framework, [24] pioneered two distinct restart strategies for extrapolation parameter: a gradient-based scheme [16, 14, 29] and a function-value-based scheme. The former, taking [29] for example, represents a computationally efficient implementation, while the latter, a function value-based restart scheme, i.e., restarting whence $E\left(x^{n}\right)>E\left(x^{n-1}\right)$, which conceptually parallels the line search techniques proposed here, provides enhanced theoretical guarantees. We also refer to [6] for other restart techniques with the estimation of Lipschitz constants based on the FISTA framework. However, the function-value-based restart scheme usually needs to be restarted frequently due to the strict energy-decreasing requirement, which leads to an increased computational cost, as observed in our numerical experiments. Crucially, we address a critical failure mode: persistent line search failures triggering an excessive restart, which degrade the algorithm to basic DCA performance. To mitigate this, we directly introduce a novel update strategy that dynamically adjusts the extrapolation parameter based on the following line search. By defining $d^{n}=\bar{x}^{n}-x^{n}$, we proceed with an aggressive non-monotone Armijo-type line search involving parameters $\rho^{k}\lambda_{\max}>0$, $k=1,2,\ldots$, $\rho\in(0,1)$, and $\lambda_{\max},\omega,\eta>0$. The aim is to find the smallest integer $k$ satisfying [15]
@@
Henceforth, denote $a(n):=k$ as the number of line searches for the $n$-th iteration and $N_{\max}$ as the maximal number of line searches. The final step size $\lambda_{n}:=\rho^{k-1}\lambda_{\max}$ obtained from the line search in (3.4) dictates the update of $x^{n}$ as follows
@@
We propose the following method to update the extrapolation parameter $\beta_{n+1}$ determined by line search (3.4). We assume that $b_{1}>0$ and $b_{2}\geq 0$ are constants close to 0 henceforth.
@@
[…]
@@
with parameter resets $\theta_{n}=\theta_{n+1}=1$ under specific conditions: fixed restart triggers resets periodically every $\bar{T}$ iterations, whereas adaptive restart activates when $y^{n}-x^{n+1},x^{n+1}-x^{n}\right>>0$.
@@
According to (3.6) for updating the extrapolation parameter $\beta_{n+1}$, we assert that there exists a positive constant $C_{\lambda}$ such that $\frac{1}{\left(1+\lambda_{n}\right)^{2}}-\beta_{n+1}^{2}>C_{\lambda}$ for $\forall n$.
@@
Setting $C_{\lambda}=\min\left(C_{\lambda}^{1},C_{\lambda}^{2}\right)$, we proved this conclusion. ∎
@@
4:Set $d^{n}=\bar{x}^{n}-x^{n}$. IF $d^{n}=0$, STOP and RETURN $x^{n+1}$. Otherwise, go to Step 8.
@@
7:Start the non-monotone line search
IF $\text{IF \quad}E\left(\bar{x}^{n}+\lambda_{n}d^{n}\right)>E\left(\bar{x}^{n}\right)-\ \lambda_{n}\left\|d^{n}\right\|^{2}+\nu_{n}.$ (3.9)
@@
4:Set $d^{n}=\bar{x}^{n}-x^{n}$. IF $d^{n}=0$, STOP and RETURN $x^{n+1}$. Otherwise, go to Step 8.
@@
7:Start the non-monotone line search
IF $\text{IF \quad}E\left(\bar{x}^{n}+\lambda_{n}d^{n}\right)>E\left(\bar{x}^{n}\right)-\eta\lambda_{n}\left\|d^{n}\right\|^{2}+\nu_{n}.$ (3.11)
@@ V LSDE non-monotone Yes Our work
Before starting the proof of convergence, we shall give the strategy [15] to choose the parameter sequence $\nu_{n}$.
@@
Let $\omega>0$ be a constant, and the sequence $\left\{\nu_{n}\right\}_{n\in N}$ defined by $\nu_{n}=\frac{\omega}{n+1}\left\|d^{n}\right\|^{2}$, for all $n\in N$. Indeed, due to $\lim_{n\rightarrow\infty}\nu_{n}=\lim_{n\rightarrow\infty}\frac{\omega}{n+1}=0$, for every $\delta>0$, there exists $n_{0}\in N$ such that $n\geq n_{0}$ implies $\frac{\omega}{n+1}\leq\delta<\eta\cdot\lambda_{\max}\cdot\rho^{N_{\max}}$. We then have $\nu_{n}\leq\delta\left\|d^{n}\right\|^{2}$. Similarly, we can define $\nu_{n}=\frac{\omega}{p\left(n\right)}\left\|d^{n}\right\|^{2}$, where $p\left(n\right)$ is a monotonically increasing function concerning $n$, here the choice of $p\left(n\right)$ depends on the rate at which we want the non-monotone term to decline.
This brings great flexibility. At the beginning of iterations, the line search (3.4) with $\nu_{n}$ is indeed non-monotone. However, it eventually becomes monotone for sufficiently large $n$ since $\nu_{n}$ converges to 00 while $N_{\max}$ is bounded.
@@
If $g_{1}+f$ is quadratic function with $\nabla(g_{1}+f)(y)=Ay-b_{0}$ and $A$ is a positive definite and linear operator $A$, the update in (3.3) can be reformulated as the following preconditioned iteration
$\bar{x}^{n}=y^{n}+\mathbb{M}^{-1}(b^{n}-Ay^{n}),$ (3.12)
@@
$\mathbb{M}=A+M,\quad b^{n}=b_{0}+\xi^{n},$
which is one time classical preconditioned iteration for solving (3.2), i.e., $Ay=b^{n}$ with initial value $y^{n}$.
@@
$y$ $=(M+A)^{-1}(My^{n}+\xi^{n}+b_{0})=(M+A)^{-1}((M+A)y^{n}+\xi^{n}+b_{0}-Ay^{n})$
$=y^{n}+(M+A)^{-1}(\xi^{n}+b_{0}-Ay^{n})=y^{n}+(M+A)^{-1}(b^{n}-Ay^{n}).$
By letting $\bar{x}^{n}=y$, the proof is finished. We emphasize that the preconditioned iteration automatically determines $M$, and no explicit expression is needed. We refer to [9, Section 2.3] for more details and [28, Proposition 3.2] for a similar idea. ∎

### v2

@@ Abstract
This paper proposes a novel proximal difference-of-convex (DC) algorithm enhanced with extrapolation and aggressive non-monotone line search for solving non-convex optimization problems. We introduce an adaptive conservative update strategy of the extrapolation parameter determined by a computationally efficient non-monotone line search. The core of our algorithm is to unite the update of the extrapolation parameter with the step size of the non-monotone line search interactively. The global convergence of the two proposed algorithms is established through the Kurdyka-Łojasiewicz properties, ensuring convergence within a preconditioned framework for linear equations. Numerical experiments on two general non-convex problems: SCAD-penalized binary classification and graph-based Ginzburg-Landau image segmentation models, demonstrate the proposed method’s high efficiency compared to existing DC algorithms in both convergence rate and solution accuracy.
@@ 1 Introduction
Here $g_{1}\left(x\right)$ and $g_{2}\left(x\right)$ are proper closed convex functions, $f$ is convex and $\nabla f\left(x\right)$ is Lipschitz continuous with constant $L$, and $X$ is a finite-dimensional Hilbert space. Such non-convex optimization formulations find broad applications in digital image processing, statistical machine learning, and related fields [1, 8, 11]. The proximal DC algorithm (DCA) with extrapolation (shortened as pDCAe henceforth) has emerged as a particularly efficient approach, successfully accelerating conventional proximal gradient methods [18, 17, 19]. This method combines the computational efficiency of proximal gradient algorithms with DC decomposition techniques, as detailed in [1, 11] for various non-convex optimizations. Our work builds upon these foundations while introducing critical enhancements to the algorithmic framework.
@@
As shown in the survey [19, Section 5], the convergence speed of DCA can be accelerated through two distinct approaches: extrapolation techniques with traditional parameterization strategies, including FISTA-based extrapolation [25, Algorithm 1] or FISTA-based restart mechanisms, e.g., [19, Algorithm 5], [29, Algorithm (3.4)] and Armijo line search methods based on decreasing energy [2, Algorithm 2], [27].
@@
Compared to the existing works, our contributions are as follows. The main contributions are a concise non-monotone line search mechanism [15] and a new adaptive update strategy for the extrapolation parameter. The line search completely determines the extrapolation parameter. Inspired by [2, 3, 15], we employ the aggressive non-monotone line search, where the starting point for line search is generated by the preconditioned and extrapolated DCA, instead of using the previous iteration [2, 3]. The aggressive non-monotone line search [15, Example 4] can not only relax stringent descent requirements, but can also obtain an extra step. The proposed framework enhances pDCAe by integrating these two accelerating methods. Specifically, the non-monotone line search strategy provides global guidance for local extrapolation parameter updates. We refer to [31, NLSA] [23, Algorithm 1] [12, Section 4] for the traditional non-monotone line search strategy. Moreover, given the limited existing work that combines these approaches, our comparative analysis demonstrates distinct advantages over existing methods. Compared to adaptive linear search methods [30, Algorithm 1], which require solving auxiliary subproblems, our approach eliminates computational overhead while preserving the core efficiency of pDCAe.
@@
Furthermore, we introduce a line search framework for the case where $g_{1}$ is non-smooth, providing a global convergence guarantee that is not covered by the line search framework [3]. In [3], only the case $g_{2}$ is non-smooth, and $g_{1}$ with Lipschitz gradient is discussed to ensure the existence of descent directions.
@@
Numerical validations confirm the improvements. In the numerical experiments in Section 5, we compare our algorithm with the algorithms that use only a single type of acceleration method. The experiments show that our algorithm has better performance. First, for iterative sequences near critical points, the integrated update strategy stabilizes convergence by controlling the extrapolation parameter $\beta$, effectively addressing the performance degradation observed in the original implementations of pDCAe. Second, the interactive combination of extrapolation and line search accelerates convergence compared to standard extrapolation and restart approaches. These empirical results, detailed in Section 5, verify that our framework maintains the computational simplicity of pDCAe while achieving better convergence across various non-convex optimization benchmarks.
@@ 2 Notation and preliminaries
Throughout this paper, we use $X$ to denote the $n$-dimensional Hilbert space with inner product $\langle\cdot,\cdot\rangle$ and Hilbert norm $\left\|\cdot\right\|$. Moreover, for a symmetric matrix $M\in\mathbb{R}^{n\times n}$, we use $\mu_{\max}\left(M\right)$ and $\mu_{\min}\left(M\right)$ to denote its largest and smallest eigenvalues respectively.
@@
$\partial h(x):=\bigg\{$ $\xi\in\mathbb{R}^{n}:\exists x_{n}\rightarrow x,h(x_{n})\rightarrow h(x),\xi_{n}\rightarrow\xi,\lim_{y\rightarrow x}\inf_{y\neq x_{n}}\frac{h(y)-h(x_{n})-\langle\xi_{n},y-x_{n}\rangle}{|y-x_{n}|}\geq 0\bigg\}.$
@@ 3 Preconditioned DCA with extrapolation determined by line search
Now let us turn to preconditioned DCA with extrapolation for (1.1). By solving the subproblem involving the difference of convex functions by incorporating a proximal term to update $\left\{\bar{x}^{n}\right\}_{n}$, the optimization task for determining $\bar{x}^{n}$ is as follows
@@
Henceforth, we assume that the symmetric positive definite matrix $M$ serves as a preconditioning weight operator, where $\langle a,b\rangle_{M}=\langle a,Mb\rangle$ for $\forall$ $a,b\in X$ and the induced norm as $\left\|x\right\|_{M}^{2}:=\left<Mx,x\right$>. Applying first-order optimality conditions to the minimization problem (3.1) yields the critical equation for determining $\bar{x}^{n}$
$\xi^{n}\in M\left(y-y^{n}\right)+\nabla f\left(y\right)+\partial g_{1}\left(y\right).$ (3.2)
@@
Here $\nabla$ denotes the gradient operator, while $\nabla f(y)\in\mathbb{R}^{n}$ represents the evaluated gradient vector at point $y$.
If $g_{1}+f$ is a quadratic function, the update $\left\{\bar{x}^{n}\right\}_{n}$ in (3.3) can be reformulated as the classical efficient preconditioned iteration, which can circumvent solving the linear equation even with mild accuracy, especially for a large-scale linear system. We will explain the preconditioning technique in Proposition 2 later.
Within the widely employed gradient descent framework, [24] pioneered two distinct restart strategies for the extrapolation parameter: a gradient-based scheme [16, 14, 29] and a function-value-based scheme. The former, taking [29] for example, represents a computationally efficient implementation, while the latter, a function value-based restart scheme, i.e., restarting when $E\left(x^{n}\right)>E\left(x^{n-1}\right)$, which conceptually parallels the line search techniques proposed here, provides enhanced theoretical guarantees. We also refer to [6] for other restart techniques with the estimation of Lipschitz constants based on the FISTA framework. However, the function-value-based restart scheme usually needs to be restarted frequently due to the strict energy-decreasing requirement, which leads to an increased computational cost, as observed in our numerical experiments. Crucially, we address a critical failure mode: persistent line search failures that trigger excessive restarts, degrading the algorithm to basic DCA performance. To mitigate this, we directly introduce a novel update strategy that dynamically adjusts the extrapolation parameter based on the following line search. By defining $d^{n}=\bar{x}^{n}-x^{n}$, we proceed with an aggressive non-monotone Armijo-type line search involving parameters $\rho^{k}\lambda_{\max}>0$, $k=1,2,\ldots$, $\rho\in(0,1)$, and $\lambda_{\max},\omega,\eta>0$. The aim is to find the smallest integer $k$ satisfying [15]
@@
Henceforth, denote $a(n):=k$ as the number of line searches for the $n$-th iteration and $N_{\max}$ as the maximal number of line searches. The final step size $\lambda_{n}:=\rho^{k-1}\lambda_{\max}$ obtained from the line search in (3.4) dictates the update of $\left\{x^{n}\right\}_{n}$ as follows
@@
We propose the following method to update the extrapolation parameter $\left\{\beta_{n+1}\right\}$ determined by line search (3.4). We assume that $b_{1}>0$ and $b_{2}\geq 0$ are constants close to 0 henceforth.
@@
$\text{Case 1: }1\leq a(n)\leq N_{\max}\text{, }\lambda_{n}=\rho^{k-1}\lambda_{\max}\text{, we set }\beta_{n+1}=\frac{1}{1+b_{1}+\lambda_{n}},$ (3.6)
$\text{Case 2: }a(n)>N_{\max}\text{, }\lambda_{n}=0\text{, we set }\beta_{n+1}=b_{2}.$
The proposed LSDE that focuses on the global energy can provide global guidance for local extrapolation parameter updates.
@@
with parameter resets $\theta_{n}=\theta_{n+1}=1$ under specific conditions: fixed restart triggers resets periodically every $\bar{T}$ iterations, whereas adaptive restart activates when $\left<y^{n}-x^{n+1},x^{n+1}-x^{n}\right>>0$.
@@
According to (3.6) for updating the extrapolation parameter $\left\{\beta_{n+1}\right\}$, we assert that there exists a positive constant $C_{\lambda}$ such that $\frac{1}{\left(1+\lambda_{n}\right)^{2}}-\beta_{n+1}^{2}>C_{\lambda}$ for $\forall n$.
@@
Setting $C_{\lambda}=\min\left(C_{\lambda}^{1},C_{\lambda}^{2}\right)$, we proved this conclusion. ∎∎
@@
4:Set $d^{n}=\bar{x}^{n}-x^{n}$. IF $d^{n}=0$, STOP and RETURN $x^{n+1}$.
@@
7:Start the non-monotone line search:
$\text{IF \quad}E\left(\bar{x}^{n}+\lambda_{n}d^{n}\right)>E\left(\bar{x}^{n}\right)-\eta\lambda_{n}\left\|d^{n}\right\|^{2}+\nu_{n}$ (3.9)
@@
4:Set $d^{n}=\bar{x}^{n}-x^{n}$. IF $d^{n}=0$, STOP and RETURN $x^{n+1}$.
@@
7:Start the non-monotone line search:
$\text{IF \quad}E\left(\bar{x}^{n}+\lambda_{n}d^{n}\right)>E\left(\bar{x}^{n}\right)-\eta\lambda_{n}\left\|d^{n}\right\|^{2}+\nu_{n}$ (3.11)
@@ V LSDE non-monotone Yes Our work
Before starting the proof of convergence, we shall give the strategy [15] to choose the parameter sequence $\{\nu_{n}\}_{n}$.
@@
Let $\omega>0$ be a constant, and the sequence $\left\{\nu_{n}\right\}_{n}$ defined by $\nu_{n}=\frac{\omega}{n+1}\left\|d^{n}\right\|^{2}$, for all $n\in N$. Indeed, due to $\lim_{n\rightarrow\infty}\nu_{n}=\lim_{n\rightarrow\infty}\frac{\omega}{n+1}=0$, for every $\delta>0$, there exists $n_{0}\in N$ such that $n\geq n_{0}$ implies $\frac{\omega}{n+1}\leq\delta<\eta\cdot\lambda_{\max}\cdot\rho^{N_{\max}}$. We then have $\nu_{n}\leq\delta\left\|d^{n}\right\|^{2}$. Similarly, we can define $\nu_{n}=\frac{\omega}{p\left(n\right)}\left\|d^{n}\right\|^{2}$, where $p\left(n\right)$ is a monotonically increasing function concerning $n$, here the choice of $p\left(n\right)$ depends on the rate at which we want the non-monotone term to decline.
This brings great flexibility. At the beginning of iterations, the line search (3.4) with $\nu_{n}$ is indeed non-monotone. However, it eventually becomes monotone for sufficiently large $n$ since $\nu_{n}$ converges to 0 while $N_{\max}$ is bounded.
@@
Suppose $(g_{1}+f)(y)=\frac{1}{2}\langle Ay,y\rangle-\langle p_{0},y\rangle+c_{0}$ with constant $c_{0}$, where $A$ is a linear, symmetric, positive definite operator and $p_{0}$ is a vector. Noting $\nabla(g_{1}+f)(y)=Ay-p_{0}$, the update in (3.3) can be reformulated as the following preconditioned iteration
$\bar{x}^{n}=y^{n}+\mathbb{M}^{-1}(p^{n}-Ay^{n}),$ (3.12)
@@
$\mathbb{M}=A+M,\quad p^{n}:=p_{0}+\xi^{n},\ \text{with}\ \xi^{n}\in\partial g_{2}\left(x^{n}\right)$
which is one time classical preconditioned iteration for solving (3.2), i.e., $Ay=p^{n}$ with initial value $y^{n}$.
@@
$y$ $=(M+A)^{-1}(My^{n}+\xi^{n}+p_{0})=(M+A)^{-1}((M+A)y^{n}+\xi^{n}+p_{0}-Ay^{n})$
$=y^{n}+(M+A)^{-1}(\xi^{n}+p_{0}-Ay^{n})=y^{n}+(M+A)^{-1}(p^{n}-Ay^{n}).$
By letting $\bar{x}^{n}=y$, the proof is finished. We emphasize that the preconditioned iteration automatically determines $M$, and no explicit expression is needed. We refer to [9, Section 2.3] for more details and [28, Proposition 3.2] for a similar idea. ∎∎